OpenStudy (teller):

A Web music store offers two versions of a popular song. The size of the standard version is 2.7 megabytes (MB). The size of the high-quality version is 4.6 MB. Yesterday, the high-quality version was downloaded twice as often as the standard version. The total size downloaded for the two versions was 2499 MB. How many downloads of the standard version were there? @Vocaloid

Vocaloid (vocaloid):

well, this is a very similar problem to the other one, so let's start off the same way. H = number of high-quality downloads S = number of standard-quality downloads H = 2*S 2.7*S + 4.6*H = 2499 do you think you can finish from here? if you get stuck, remember the other problem that we just did

OpenStudy (teller):

I don't have the other problem since it got rid of it so do I take 2.7(3-S)+4.6=2499?

Vocaloid (vocaloid):

I'm not sure where you're getting 3 from

OpenStudy (teller):

i mean 2 sorry

Vocaloid (vocaloid):

remember, H = 2*S means that you can replace "H" with "2S" whenever you see it

Vocaloid (vocaloid):

take another look if you are still confused http://openstudy.com/study#/updates/55d52a97e4b0554d6272fa4c

OpenStudy (teller):

so2.7s+4.6(2-s)=2499?

Vocaloid (vocaloid):

almost, there shouldn't be a minus sign... 2.7S + 4.6(2*S) = 2499, agreed?

OpenStudy (teller):

alright so if im correct i'd take 4.6 and times it by 2?

Vocaloid (vocaloid):

right, and that gives us...?

OpenStudy (teller):

9.2?

Vocaloid (vocaloid):

don't forget the S 4.2(2*S) = 9.2S

Vocaloid (vocaloid):

so that leaves us with 2.7*S + 9.2*S = 2499 can you try solving for S now?

OpenStudy (teller):

would it be 210?

Vocaloid (vocaloid):

yup, so that's our answer

OpenStudy (teller):

thank you

Vocaloid (vocaloid):

S = number of standard downloads = 210